3x - 4y + 3 = 0
6x - 8y +7 = 0
a. Find the distance between the two lines.
b. Find the equation of the perpendicular line passing through. (-6,4)
c. Determine the distance of 2 given equations to point (-6.4)
1
Expert's answer
2020-05-08T20:09:35-0400

Let two lines be defined by equations "A_1x+B_1y+C_1=0" and "A_2x+B_2y+C_2=0" . These lines are parallel if "A_2B_1-A_1B_2=0." For our lines "6\cdot(-4)-3\cdot(-8)=0" , so they are parallel. Let us rewrite the equation of the second line as

"3x-4y+3.5=0."

Therefore, our lines are defined by equations

"A_1x+B_1y+C_1=0, \;\; A_1x+B_1y+C=0."

a) Now the distance can be calculated as (see https://en.wikipedia.org/wiki/Distance_between_two_straight_lines)

"d= \dfrac{|C_1-C|}{\sqrt{A^2+B^2}} = \dfrac{|3-3.5|}{\sqrt{3^2+4^2}} = 0.1."


b) Let the line be defined by equation "A_0x+B_0y+C_0=0." This line is perpendicular to "A_1x+B_1y+C_1=0" if "A_0A_1+B_0B_1 = 0" . Let us choose "A_0 = 4, \; B_0= 3."

We know that "A_0\cdot(-6)+B_0\cdot4+C_0=0." Therefore, "C_0=12."

So the equation of line is "4x+3y+12=0."


c) Distance from point "(x_0,y_0)" to line "Ax+By+C=0" (see https://en.wikipedia.org/wiki/Distance_from_a_point_to_a_line)

"d=\dfrac{|Ax_0+By_0+C|}{\sqrt{A^2+B^2}}" .

For the first line it is

"d_1=\dfrac{|3\cdot(-6)-4\cdot4+3|}{\sqrt{3^2+4^2}} = \dfrac{31}{5} =6.2 ."

For the second line it is

"d_2=\dfrac{|3\cdot(-6)-4\cdot4+3.5|}{\sqrt{3^2+4^2}} = \dfrac{30.5}{5} = 6.1."